OpenStudy (anonymous):

A vertical pole 27 feet tall stands on a hillside that makes an angle of 19o with the horizontal. Determine the approximate length of cable that would be needed to reach from the top of the pole to a point 80 feet downhill from the base of the pole. Round answer to two decimal places.

OpenStudy (anonymous):

these are the options: 92.39 feet B. 100.82 feet C. 84.02 feet D. 89.62 feet E. 75.65 feet

OpenStudy (anonymous):

92.39 feet We will need to use the law of cosines. I think the tough part is the meaning of 19 degrees with the horizon; this means the angle is 109 = 90 + 19 Using the law of cosines we have c^2 = 80^2 + 27^2 - 2(80)(27)cos(109) and taking the square root of both sides we have c = 92.39 feet

OpenStudy (anonymous):

Thank you so much! :)